1/2

Bob Dylon ,  Bob Dylon ,  Bob Dylon ,

Blowin’ in the Wind Kiom da mono elspezas la hom’ Die Antwort weiß ganz allein der Wind

translated by Manfred Retzlaff

translated by Hans Bradtke

How many roads must a man walk

down Before you call him a man? How many seas must the white do-

ve sail Before she sleeps in the sand?

Kiom da mono elspezas la

hom’ Por la armilar’ sur la ter’?

- Kaj kiom da vortoj dira-

tas de li? Kaj ne malpliiĝas mizer’!

Wie viele Straßen auf dieser

Welt sind Straßen voll Tränen und

Leid? Wie viele Meere auf dieser Welt sind Meere der Einsamkeit?

Yes, and how many times must the

cannonballs fly Before they are forever banned? The answer, my friend, is blowin’

in the wind The answer is blowin’ in the wind Ĉu devos okazi la lasta mi-

lit’ Kun neimagebla mizer’? Respondon, amik’, ja sci-

as nur la vent’, Repondon ja scias nur la

vent’.

Wie viele Mütter sind lang schon

allein und warten und warten noch

heut? Die Antwort, mein Freund, sie

will gefunden sein, Die Antwort muss unser Leben

sein.

Yes, and how many years can a

mountain exist Before it is washed to the sea? Yes, and how many years can some

people exist Before they’re allowed to be free?

Kiom da stratoj sur tiu ĉi

ter’ Kondukas al senesperec’?

- Kaj tiom da maroj en tiu

ĉi mond’ Plenplenas de plor’ kaj

tristec’? - Wie viele Menschen sind heut

noch nicht frei und würden es so gerne sein? Wie viele Kinder gehn abends

zur Ruh und schlafen vor Hunger nicht

ein?

Yes, and how many times can a

man turn his head And pretend that he just doesn’t

see? The answer, my friend, is blowin’

in the wind The answer is blowin’ in the wind Kaj kiom da inoj atendas,

ho ve, Kaj restas en tia solec’? Respondon, amik’, ja sci-

as nur la vent’, Repondon ja scias nur la

vent’.

Wie viele Tränen erflehen bei

Nacht: Wann wird es für uns anders

sein? Die Antwort, mein Freund, sie

will gefunden sein, Die Antwort muss unser Leben

sein.

…  …  …

2/2

Yes, and how many times must a

man look up Before he can see the sky? Yes, and how many ears must one

man have Before he can hear people cry?

Kiom da homoj en malli-

berec’ Suferas nur pro opini’? - Kaj kiom da homaj infa-

noj, ho jes, Malsatas sur ter tiu ĉi?

Wie viele Berge von Geld gibt

man aus für Bomben, Raketen und Tod? Wie gro0e Worte macht heut

mancher Mann und lindert damit keine Not?

Yes, and how many deaths will it

take ’til he knows That too many people have died? The answer, my friend, is blowin’

in the wind The answer is blowin’ in the wind.

Kaj kial la revoj sur pac’

sur al ter’ Ne efektiviĝas sur ĝi? Respondon, amik’, ja sci-

as nur la vent’, Repondon ja scias nur la

vent’.

Wie großes Unheil muss erst

noch geschehn, damit sich die Menschheit be-

sinnt? Die Antwort, mein Freund, sie

will gefunden sein, Die Antwort muss unser Leben

sein.

Author of this English poem is Bob Dylon (civila nomo: Robert Allen Zimmermann, ∗1941-05-24), published 1962-04.

Arg-720-1446 (2012-12-11 09:07:24)

Pri la aŭtoro, vidu la retejon http: // www. songlexikon. de/ songs/ blowinwind , pri la teksto la retejon http: // www. songtexte. com/ songtext/ b b-dylan/ blowin-in-the-wind-2bd5e8d2. html . La melodio estas aŭdebla en http: // www. youtube. com/ watch? v= tGO8snK98l0 kaj en http: // www. myvideo. de/ watch/ 3594984/ BOB_DYLAN_BLOWIN_IN_ THE_WIND .

Translation of the English poem “Blowin’ in the Wind” by Bob Dylon (civila nomo: Robert Allen Zimmermann, ∗1941-05-24) into Esperanto by Manfred Retzlaff (∗1938-11-04) on 1984-04-29.

MR-148-1 / Arg-720-1447 (2012-12-11 09:05:35)

La melodio de tiu ĉi kanto, kiun verkis la aŭtoro Bob Dylon (civila nomo: Robert Allen Zimmermann) ankaŭ, aŭdeblas en http: // www. youtube. com/ watch? v= tGO8snK98l0 . Pri la muziknotoj vidu: http: // de. scribd. com/ doc/ 9589786/ Bob-Dylan-Blowin-in-the-wind-notes .

Translation of the English poem “Blowin’ in the Wind” by Bob Dylon (civila nomo: Robert Allen Zimmermann, ∗1941-05-24) into German by Hans Bradtke (∗1920-07-21 – †1997-05-12).

Arg-720-1448 (2012-12-11 09:13:17)

Siehe  auch:  http: // www.

sjd-falken. de/ lieder/ die antwort_weiss_ganz_ allein_der_wind. shtml . Pri la tradukinto vidu: http: // www. discogs. com/ artist/ Hans+ Bradtke? anv= H. +Bradtke&filter_anv= 1 kaj la vikipediejon http: // de. wikipedia. org/ wiki/ Hans_Bradtke .

http://www.poezio.net/version?poem-id=720&version-id=1446,1447,1448